OpenStudy (marigirl):

Having a blank moment.. please help me write square root of u on the top line

OpenStudy (marigirl):

\[\frac{ 2u-1 }{ \sqrt{u} }\]

OpenStudy (marigirl):

does it become (2u-1)(u)^(-0.5)

OpenStudy (compassionate):

Do you want o solve for u?

OpenStudy (marigirl):

i want to bring square root of u to the top

OpenStudy (marigirl):

\[\frac{ 2u-1 }{ \sqrt{u} }=2u-2*x ^{\frac{ -1 }{ 2 }}\]

OpenStudy (marigirl):

is that correct?

OpenStudy (marigirl):

i mean 2u-1

OpenStudy (compassionate):

\[\frac{ (2u-1^{(0.05u)} }{ 1 }\]

OpenStudy (marigirl):

omg i cant do that....because its a relationship at the top!!!!!! opps

OpenStudy (anonymous):

(2u−1(0.05u)1

OpenStudy (marigirl):

\[\frac{ 2u-1 }{ \sqrt{u} }=\frac{ 2u }{ \sqrt{u} }-\frac{ 1 }{ \sqrt{u} }\]

zepdrix (zepdrix):

If you multiply \(\rm \sqrt u\) by \(\rm \sqrt u\) will get rid of the square root down there :) ya?

zepdrix (zepdrix):

\[\large\rm \frac{(2u-1)}{\sqrt{u}}\cdot\color{royalblue}{\frac{\sqrt u}{\sqrt u}}\]Gotta do this fancy business

OpenStudy (anonymous):

this is ok because your effectively multiplying by 1